# Singularity-free cosmology from interactions in the dark sector

### v1

@@ IV.1.1 The repellor-saddle separatrix and the $a^{\prime\prime}=0$ curve do not intersect
The first subcase when $dS_{1+}$ is a spiral repellor is shown in Fig. 2.555We plot $Z$ on the horizontal axis and $X$ on the vertical axis so the 2-D phase spaces correspond more clearly to the 3-D phase spaces, where $X$ is plotted on the vertical axis and $Z$ is plotted on the depth axis. Here, the $a^{\prime\prime}=0$ curve and repellor-saddle separatrix do not intersect, and not all trajectories have a decelerated period. Those trajectories that do not intersect the zero acceleration curve correspond to first integral surfaces in the $X$-$Y$-$Z$ phase space that have no Einstein fixed points. Some of these trajectories expand to the left of the two separatrices and evolve towards $S_{2+}$, and some evolve to the right of the two separatrices and asymptotically expand towards the late-time attractor $dS_{2+}$. Both of these sets of trajectories are always accelerating, and so are not of interest. One trajectory in the phase space touches the $a^{\prime\prime}=0$ curve, and corresponds to a first integral surface in 3-D that has one Einstein fixed point. This trajectory evolves toward $dS_{2+}$, and will have zero acceleration where it touches the zero acceleration curve, but does not have a decelerated period so is not qualitatively interesting. The rest of the trajectories intersect the $a^{\prime\prime}=0$ curve twice, which correspond to first integral surfaces in 3-D with two Einstein fixed points. These trajectories initially accelerate, then cross below the $a^{\prime\prime}=0$ curve and decelerate, and finally cross back above $a^{\prime\prime}=0$ and accelerate as they asymptotically approach the late-time attractor $dS_{2+}$. These trajectories are of interest as qualitatively they have a decelerated period where large-scale structure could form, followed by a late-time accelerated expansion that tends toward a cosmological constant.
@@
Figure 2: The $Z$-$X$ phase space referred to in Sec. IV.1.1, where parameters are set to $w_{x}=-0.65$ and $q=1.5$. Fig 2(a) shows the full phase space. Trajectories that evolve to the left of the two separatrices expand to the critical point $S_{2+}$ at $(1,1)$ representing a singularity, and those that evolve to the right expand towards the attractor fixed point $dS_{2+}$ (magenta). The zero acceleration curve does not intersect the repellor-saddle separatrix, therefore all trajectories that evolve to the left of the separatrices, and those that evolve to the right of the separatrices that do not intersect the $a^{\prime\prime}=0$ curve, always accelerate. Trajectories that intersect the $a^{\prime\prime}=0$ curve twice have a decelerated period, followed by a late-time acceleration. The trajectory that touches the zero acceleration curve reaches a point where there is zero acceleration, but accelerates otherwise. Fig 2(b) is an enlargement of panel (a) around the origin, showing the behavior of the trajectories that evolve to $dS_{2+}$. Those that evolve to $dS_{2+}$ from above the green $X^{\prime}=0$ curve have non-phantom behavior, with decreasing dark energy density. Some trajectories have late-time phantom behavior, and cross below the green $X^{\prime}=0$ curve and approach $dS_{2+}$ with increasing dark energy density. All trajectories accelerate as they approach $dS_{2+}$, as they are to the left of the red $a^{\prime\prime}=0$ curve.
@@ IV.1.2 The repellor-saddle separatrix and the $a^{\prime\prime}=0$ curve touch
Fig. 3 shows the subcase where the repellor-saddle separatrix and zero acceleration curve touch. The repellor-saddle separatrix itself corresponds to a first integral surface in the $X$-$Y$-$Z$ phase space with one Einstein fixed point. This trajectory initially accelerates, reaches a point with zero acceleration and then accelerates again, and tends towards the saddle fixed point $dS_{3+}$. As this trajectory has no decelerated period, it is not of interest. Trajectories that evolve to the left of the two separatrices correspond to first integral surfaces in 3-D that have zero Einstein fixed points, and never decelerate. These trajectories also evolve to a singularity, therefore they are not of interest. The rest of the trajectories evolve to the right of the separatrices, and intersect the $a^{\prime\prime}=0$ curve twice, meaning they have a decelerated period. They also all tend towards $dS_{2+}$ at late-times, and are therefore the trajectories of interest.
Figure 3: The $Z$-$X$ phase space referred to in Sec. IV.1.2. Here the parameters are set to $w_{x}=-0.72$ and $q=1.5$. Trajectories that evolve to the left of the two separatrices expand to the critical point $S_{2+}$ at $(1,1)$ representing a singularity, and those that evolve to the right expand towards the attractor fixed point $dS_{2+}$ (magenta). In this case, the zero acceleration curve touches the repellor-saddle separatrix. Trajectories that evolve to the left of the separatrices therefore always accelerate, and those that evolve to the right all have a decelerated period. The repellor-saddle separatrix is a trajectory that has a point of zero acceleration, but accelerates otherwise.
@@ IV.1.3 The repellor-saddle separatrix and the $a^{\prime\prime}=0$ curve intersect
The final subcase is where the repellor-saddle separatrix intersects the $a^{\prime\prime}=0$ curve twice, as shown in Fig. 4. The trajectories that evolve to the left of the two separatrices all tend toward $S_{2+}$; those that intersect the $a^{\prime\prime}=0$ curve twice have a period of deceleration, and those that do not intersect $a^{\prime\prime}=0$ always accelerate. Regardless, these trajectories are not interesting with respect to our analysis as they evolve towards a singularity. Trajectories to the right of the separatrices all have a decelerated period followed by a late-time acceleration, where they tend toward $dS_{2+}$. These trajectories are all of interest, therefore this is the subcase that we focus our analysis in 3-D on in the following section.
Figure 4: The $Z$-$X$ phase space referred to in Sec. IV.1.3. Here, the parameters are set to $w_{x}=-0.9$ and $q=1.5$. Trajectories that evolve to the left of the two separatrices expand to the critical point $S_{2+}$ at $(1,1)$ representing a singularity, and those that evolve to the right expand towards the attractor fixed point $dS_{2+}$ (magenta). In this case, the zero acceleration curve intersects the repellor-saddle separatrix twice, therefore trajectories that evolve to the right of the separatrices all have a decelerated phase followed by late-time acceleration.
IV.2 3-D Phase Space for Viable Models
@@
[…]
$dS_{3\pm}$ $1+w_{x}$$\pm\sqrt{\frac{(1+w_{x})}{3}}$ 00
Table 6: The fixed points of the full 3-D system in Eq.s (9) - (11) with $\epsilon=-1$. $E$ denotes a static Einstein universe ($y^{\prime}=y=0$) and $dS_{\pm}$ an expanding (+) or contracting (-) de-Sitter universe ($x^{\prime}=z^{\prime}=0$).
@@
$dS_{5\pm}$ $\frac{\mathcal{R}}{1+\mathcal{R}}$$\pm 1$ 00
$dS_{6\pm}$ $\frac{1+w_{x}}{2+w_{x}}$$\pm 1$ 00
$S_{1\pm}$ $1$$\pm 1$ 00
$S_{2\pm}$ $1$$\pm 1$ $1$
$S_{3\pm}$ 00 ±$\pm 1$ $1$
Table 7: The critical points that arise at $y\rightarrow\infty$ from compactification of the full system in Eq.s (20) - (22), with $\epsilon=-1$. $dS_{\pm}$ denotes an expanding (+) or contracting (-) de-Sitter universe ($X^{\prime}=Z^{\prime}=0$), and $S_{\pm}$ denotes a singularity with infinite expansion (+) or contraction (-).
@@
Table 8: The colors of the fixed points and critical points in the $X$-$Y$-$Z$ phase spaces. $E$ denotes a static Einstein universe and $dS_{\pm}$ an expanding (+) or contracting (-) de-Sitter universe ($X^{\prime}=Z^{\prime}=0$). $S_{\pm}$ denotes a singularity with infinite expansion (+) or contraction (-).
@@
Closed Friedmann Separatrix (CFS) Black curve
@@
Table 9: The colors of different features in the 3-D $X$-$Y$-$Z$ phase spaces.
@@
In this subsection, we focus our analysis on the set of trajectories that evolve to the right of the separatrices in Fig. 4. The fixed points of the full system are shown in Table 6 and the critical points are shown in Table 7. Their color in the 3-D plots are shown in Table 8, which correlates to the 2-D phase spaces. The color of the different features in the 3-D phase spaces are given in Table 9. We plot on a constant $Z_{0}$ and $X_{0}$ surface in 3-D, meaning the $Z$-$X$ plane in 3-D at any $Y$ looks the same as the 2-D phase space.
@@
Fig.s 5 and 6 show a first integral surface in the $X$-$Y$-$Z$ phase space with two Einstein fixed points, which corresponds to a trajectory in Fig. 4 that evolves to the right of the separatrices and intersects the $a^{\prime\prime}=0$ curve twice. In this case, one Einstein fixed point is a saddle and one is a center. A separatrix exists through the saddle Einstein fixed point, which we call the Closed Friedmann Separatrix (CFS) Ananda and Bruni (2006a), which separates different types of trajectories in the 3-D phase space. Fig. 5 shows the trajectories with positive spatial curvature only, and Fig. 6 shows the trajectories corresponding to expanding flat and open models.
@@
Figure 5: The $X$-$Y$-$Z$ phase space corresponding to Fig. 4, where $w_{x}=-0.9$ and $q=1.5$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.1$ and $Z_{0}=0.5$, which corresponds to a trajectory in Fig. 4 that evolves to the right of the separatrices, expanding to $dS_{2+}$. In this figure, only the trajectories with positive spatial curvature are plotted. Two Einstein fixed points (cyan) exist at $Y=0$ on the zero acceleration surface. Trajectories inside the CFS either bounce once, contracting from $dS_{2-}$ and expanding to $dS_{2+}$, or are cyclic around an Einstein fixed point. The bouncing models inside the CFS always accelerate, and the cyclic models cross the zero acceleration surface once during expansion, meaning they have a decelerated phase but no late-time acceleration. Trajectories outside the CFS bounce once between $dS_{2-}$ and $dS_{2+}$, and during expansion cross the zero acceleration surface twice, meaning they have a decelerated phase and a late-time acceleration.
@@
The trajectories in Fig. 5 all have a bounce. The trajectories that evolve outside the CFS, as well as a subset of those that evolve inside the CFS, bounce once. These trajectories contract from $dS_{2-}$, go through a non-singular bounce, and then expand towards $dS_{2+}$. There is also a subset of cyclic models within the CFS. These cycle around the Einstein fixed point which is a center, and repeatedly expand, reach a turnaround, contract and bounce. The trajectories that evolve inside the CFS are not of physical interest, as qualitatively they do not match the observed Universe. The cyclic models accelerate through the bounce, and then have a decelerated expanding period, however they never have a late-time accelerated expansion  [^6] For an example of cyclic models that have a decelerated phase followed by a late-time accelerated expansion, see Bruni et al. (2022).. The bouncing models that evolve inside the CFS always accelerate. The trajectories that bounce once outside of the CFS are the models of interest, as they have an accelerated era evolving toward a high energy quasi-de Sitter phase with a bounce, becoming flatter as they start to expand, then have a decelerated expanding phase followed by a final period of acceleration as they evolve toward the low energy cosmological constant represented by $dS_{2+}$.
@@
Figure 6: The expanding flat (green) and open (blue) trajectories in the $X$-$Y$-$Z$ phase space corresponding to Fig. 4, where $w_{x}=-0.9$ and $q=1.5$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.1$ and $Z_{0}=0.5$, which corresponds to a trajectory in Fig. 4 that evolves to the right of the separatrices, expanding to $dS_{2+}$. The flat and open trajectories emerge from $dS_{1+}$ and $dS_{4+}$, respectively, and both cross the zero acceleration surface twice, meaning they initially accelerate, then have a period of deceleration, then have an accelerated phase as they expand toward $dS_{2+}$ at low energy.
@@
The expanding open and flat trajectories in Fig. 6 are all qualitatively interesting. These models emerge from $dS_{1+}$ and $dS_{4+}$, respectively, and initially accelerate. They then have a decelerated period as they cross below the $a^{\prime\prime}=0$ surface, and have a final late-time acceleration as they asymptotically tend toward $dS_{2+}$. We note that although the system can violate the NEC, only closed models can have a bounce as we only consider the region of phase space where we have positive energy densities ($10<X<$ 1, $10<Z<$ 1). Therefore, $H\neq 0$ when $k\leq 0$.

### v2

@@ IV.1.1 The repellor-saddle separatrix and the $a^{\prime\prime}=0$ curve do not intersect
The first subcase when $dS_{1+}$ is a spiral repellor is shown in Fig. 2.555We plot $Z$ on the horizontal axis and $X$ on the vertical axis so the two-dimensional phase spaces correspond more clearly to the three-dimensional phase spaces, where $X$ is plotted on the vertical axis and $Z$ is plotted on the depth axis. Here, the $a^{\prime\prime}=0$ curve and repellor-saddle separatrix do not intersect, and not all trajectories have a decelerated period. Those trajectories that do not intersect the zero acceleration curve correspond to first integral surfaces in the $X$-$Y$-$Z$ phase space that have no Einstein fixed points. Some of these trajectories expand to the left of the two separatrices and evolve towards $S_{2+}$, and some evolve to the right of the two separatrices and asymptotically expand towards the late-time attractor $dS_{2+}$. Both of these sets of trajectories are always accelerating, and so are not of interest. One trajectory in the phase space touches the $a^{\prime\prime}=0$ curve, and corresponds to a first integral surface in three dimensions that has one Einstein fixed point. This trajectory evolves toward $dS_{2+}$, and will have zero acceleration where it touches the zero acceleration curve, but does not have a decelerated period so is not qualitatively interesting. The rest of the trajectories intersect the $a^{\prime\prime}=0$ curve twice, which correspond to first integral surfaces in three dimensions with two Einstein fixed points. These trajectories initially accelerate, then cross below the $a^{\prime\prime}=0$ curve and decelerate, and finally cross back above $a^{\prime\prime}=0$ and accelerate as they asymptotically approach the late-time attractor $dS_{2+}$. These trajectories are of interest as qualitatively they have a decelerated period where large-scale structure could form, followed by a late-time accelerated expansion that tends toward a cosmological constant.
@@
Figure 2: $Z$-$X$ phase space referred to in Sec. IV.1.1, where parameters are set to $w_{x}=-0.65$ and $q=1.5$. (2(a)) shows the full phase space. Trajectories that evolve to the left of the two separatrices expand to the critical point $S_{2+}$ at $(1,1)$ representing a singularity, and those that evolve to the right expand towards the attractor fixed point $dS_{2+}$ (magenta). The zero acceleration curve does not intersect the repellor-saddle separatrix, therefore all trajectories that evolve to the left of the separatrices, and those that evolve to the right of the separatrices that do not intersect the $a^{\prime\prime}=0$ curve, always accelerate. Trajectories that intersect the $a^{\prime\prime}=0$ curve twice have a decelerated period, followed by a late-time acceleration. The trajectory that touches the zero acceleration curve reaches a point where there is zero acceleration, but accelerates otherwise. (2(b)) is an enlargement of panel (2(a)) around the origin, showing the behavior of the trajectories that evolve to $dS_{2+}$. Those that evolve to $dS_{2+}$ from above the green $X^{\prime}=0$ curve have nonphantom behavior, with decreasing dark energy density. Some trajectories have late-time phantom behavior, and cross below the green $X^{\prime}=0$ curve and approach $dS_{2+}$ with increasing dark energy density. All trajectories accelerate as they approach $dS_{2+}$, as they are to the left of the red $a^{\prime\prime}=0$ curve.
@@ IV.1.2 The repellor-saddle separatrix and the $a^{\prime\prime}=0$ curve touch
Figure 3 shows the subcase where the repellor-saddle separatrix and zero acceleration curve touch. The repellor-saddle separatrix itself corresponds to a first integral surface in the $X$-$Y$-$Z$ phase space with one Einstein fixed point. This trajectory initially accelerates, reaches a point with zero acceleration and then accelerates again, and tends towards the saddle fixed point $dS_{3+}$. As this trajectory has no decelerated period, it is not of interest. Trajectories that evolve to the left of the two separatrices correspond to first integral surfaces in three dimensions that have zero Einstein fixed points, and never decelerate. These trajectories also evolve to a singularity, therefore they are not of interest. The rest of the trajectories evolve to the right of the separatrices, and intersect the $a^{\prime\prime}=0$ curve twice, meaning they have a decelerated period. They also all tend towards $dS_{2+}$ at late times, and are therefore the trajectories of interest.
Figure 3: $Z$-$X$ phase space referred to in Sec. IV.1.2. Here the parameters are set to $w_{x}=-0.72$ and $q=1.5$. Trajectories that evolve to the left of the two separatrices expand to the critical point $S_{2+}$ at $(1,1)$ representing a singularity, and those that evolve to the right expand towards the attractor fixed point $dS_{2+}$ (magenta). In this case, the zero acceleration curve touches the repellor-saddle separatrix. Trajectories that evolve to the left of the separatrices therefore always accelerate, and those that evolve to the right all have a decelerated period. The repellor-saddle separatrix is a trajectory that has a point of zero acceleration, but accelerates otherwise.
@@ IV.1.3 The repellor-saddle separatrix and the $a^{\prime\prime}=0$ curve intersect
The final subcase is where the repellor-saddle separatrix intersects the $a^{\prime\prime}=0$ curve twice, as shown in Fig. 4. The trajectories that evolve to the left of the two separatrices all tend toward $S_{2+}$; those that intersect the $a^{\prime\prime}=0$ curve twice have a period of deceleration, and those that do not intersect $a^{\prime\prime}=0$ always accelerate. Regardless, these trajectories are not interesting with respect to our analysis as they evolve towards a singularity. Trajectories to the right of the separatrices all have a decelerated period followed by a late-time acceleration, where they tend toward $dS_{2+}$. These trajectories are all of interest, therefore this is the subcase that we focus our analysis in three dimensions on in the following section.
Figure 4: $Z$-$X$ phase space referred to in Sec. IV.1.3. Here, the parameters are set to $w_{x}=-0.9$ and $q=1.5$. Trajectories that evolve to the left of the two separatrices expand to the critical point $S_{2+}$ at $(1,1)$ representing a singularity, and those that evolve to the right expand towards the attractor fixed point $dS_{2+}$ (magenta). In this case, the zero acceleration curve intersects the repellor-saddle separatrix twice, therefore trajectories that evolve to the right of the separatrices all have a decelerated phase followed by late-time acceleration.
IV.2 Three-dimensional phase space for viable models
@@
$dS_{2\pm}$ $\mathcal{R}$ $\pm\sqrt{\frac{\mathcal{R}}{3}}$ 0
$dS_{3\pm}$ $1+w_{x}$ $\pm\sqrt{\frac{(1+w_{x})}{3}}$ 0
Table 6: Fixed points of the full three-dimensional system in Eqs. (10) – (12) with $\epsilon=-1$. $E$ denotes a static Einstein universe ($y^{\prime}=y=0$) and $dS_{\pm}$ an expanding (+) or contracting (-) de Sitter universe ($x^{\prime}=z^{\prime}=0$).
@@
$dS_{5\pm}$ $\frac{\mathcal{R}}{1+\mathcal{R}}$ $\pm 1$ 0
$dS_{6\pm}$ $\frac{1+w_{x}}{2+w_{x}}$ $\pm 1$ 0
$S_{1\pm}$ 11 $\pm 1$ 0
$S_{2\pm}$ 11 $\pm 1$ 11
$S_{3\pm}$ 0 $\pm 1$ 11
Table 7: Critical points that arise at $y\rightarrow\infty$ from compactification of the full system in Eqs. (21) – (23), with $\epsilon=-1$. $dS_{\pm}$ denotes an expanding (+) or contracting (-) de Sitter universe ($X^{\prime}=Z^{\prime}=0$), and $S_{\pm}$ denotes a singularity with infinite expansion (+) or contraction (-).
@@
Table 8: Colors of the fixed points and critical points in the $X$-$Y$-$Z$ phase spaces. $E$ denotes a static Einstein universe and $dS_{\pm}$ an expanding (+) or contracting (-) de Sitter universe ($X^{\prime}=Z^{\prime}=0$). $S_{\pm}$ denotes a singularity with infinite expansion (+) or contraction (-).
@@
Closed Friedmann separatrix Black curve
@@
Table 9: Colors of different features in the three-dimensional $X$-$Y$-$Z$ phase spaces.
@@
In this subsection, we focus our analysis on the set of trajectories that evolve to the right of the separatrices in Fig. 4. The fixed points of the full system are shown in Table 6 and the critical points are shown in Table 7. Their color in the three-dimensional plots are shown in Table 8, which correlates to the two-dimensional phase spaces. The color of the different features in the three-dimensional phase spaces are given in Table 9. We plot on a constant $Z_{0}$ and $X_{0}$ surface in three dimensions, meaning the $Z$-$X$ plane in three dimensions at any $Y$ looks the same as the two-dimensional phase space.
@@
Figures 5 and 6 show a first integral surface in the $X$-$Y$-$Z$ phase space with two Einstein fixed points, which corresponds to a trajectory in Fig. 4 that evolves to the right of the separatrices and intersects the $a^{\prime\prime}=0$ curve twice. In this case, one Einstein fixed point is a saddle and one is a center. A separatrix exists through the saddle Einstein fixed point, which we call the closed Friedmann separatrix (CFS) Ananda and Bruni (2006a), which separates different types of trajectories in the three-dimensional phase space. Figure 5 shows the trajectories with positive spatial curvature only, and Fig. 6 shows the trajectories corresponding to expanding flat and open models.
@@
Figure 5: $X$-$Y$-$Z$ phase space corresponding to Fig. 4, where $w_{x}=-0.9$ and $q=1.5$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.1$ and $Z_{0}=0.5$, which corresponds to a trajectory in Fig. 4 that evolves to the right of the separatrices, expanding to $dS_{2+}$. In this figure, only the trajectories with positive spatial curvature are plotted. Two Einstein fixed points (cyan) exist at $Y=0$ on the zero acceleration surface. Trajectories inside the CFS either bounce once, contracting from $dS_{2-}$ and expanding to $dS_{2+}$, or are cyclic around an Einstein fixed point. The bouncing models inside the CFS always accelerate, and the cyclic models cross the zero acceleration surface once during expansion, meaning they have a decelerated phase but no late-time acceleration. Trajectories outside the CFS bounce once between $dS_{2-}$ and $dS_{2+}$, and during expansion cross the zero acceleration surface twice, meaning they have a decelerated phase and a late-time acceleration.
@@
The trajectories in Fig. 5 all have a bounce. The trajectories that evolve outside the CFS, as well as a subset of those that evolve inside the CFS, bounce once. These trajectories contract from $dS_{2-}$, go through a nonsingular bounce, and then expand towards $dS_{2+}$. There is also a subset of cyclic models within the CFS. These cycle around the Einstein fixed point which is a center, and repeatedly expand, reach a turnaround, contract and bounce. The trajectories that evolve inside the CFS are not of physical interest, as qualitatively they do not match the observed Universe. The cyclic models accelerate through the bounce, and then have a decelerated expanding period; however, they never have a late-time accelerated expansion  [^6] For an example of cyclic models that have a decelerated phase followed by a late-time accelerated expansion, see Ref. Bruni et al. (2022).. The bouncing models that evolve inside the CFS always accelerate. The trajectories that bounce once outside of the CFS are the models of interest, as they have an accelerated era evolving toward a high energy quasi-de Sitter phase with a bounce, becoming flatter as they start to expand, then have a decelerated expanding phase followed by a final period of acceleration as they evolve toward the low energy cosmological constant represented by $dS_{2+}$.
@@
Figure 6: Expanding flat (green) and open (blue) trajectories in the $X$-$Y$-$Z$ phase space corresponding to Fig. 4, where $w_{x}=-0.9$ and $q=1.5$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.1$ and $Z_{0}=0.5$, which corresponds to a trajectory in Fig. 4 that evolves to the right of the separatrices, expanding to $dS_{2+}$. The flat and open trajectories emerge from $dS_{1+}$ and $dS_{4+}$, respectively, and both cross the zero acceleration surface twice, meaning they initially accelerate, then have a period of deceleration, then have an accelerated phase as they expand toward $dS_{2+}$ at low energy.
@@
The expanding open and flat trajectories in Fig. 6 are all qualitatively interesting. These models emerge from $dS_{1+}$ and $dS_{4+}$, respectively, and initially accelerate. They then have a decelerated period as they cross below the $a^{\prime\prime}=0$ surface, and have a final late-time acceleration as they asymptotically tend toward $dS_{2+}$. We note that although the system can violate the NEC, only closed models can have a bounce as we only consider the region of phase space where we have positive energy densities ($0<X<1$, $0<Z<1$). Therefore, $H\neq 0$ when $k\leq 0$.